Determine the exact value of each trigonometric ratio

a) sin 75 degree
b) tan 5pie over 12
c) sin -pie over 12
1
Expert's answer
2020-10-29T15:33:22-0400

a)sin(75o) = sin(30o + 45o) = sin(30o)*cos(45o) + sin(45o)*cos(30o) = ("1\/2")*("\sqrt2/2" ) + "(\sqrt2/2)*\sqrt3/2)" = "(\sqrt2/2)*((1+\sqrt3)/2) = 0.9659"


b)tan(5*"\pi"/12) = tan(3*"\pi"/12 + 2*"\pi"/12) = tan("\pi"/4 + "\pi"/6) = (tan("\pi"/4) + tan("\pi"/6))/(1-tan("\pi"/4)*tan("\pi"/6)) = ("1 +""\sqrt3/3"")/(1-1*\sqrt3/3)" = 3.732


c)sin(-"\pi"/12) = sin(2"\pi"/12 - 3"\pi"/12) = sin(2"\pi"/12)*cos(3"\pi"/12) - sin(3"\pi"/12)*cos(2"\pi"/12) =

sin("\pi"/6)*cos("\pi"/4) - sin("\pi"/4)*cos("\pi"/6) = "(1/2)*(\sqrt2/2) - (\sqrt2/2)*(\sqrt3/2) = 
(\sqrt2/2)*(1-\sqrt3) = -0.5176"